8. A diesel engine that develops 30 kW at 3600 rpm delivers its power through a 4:1 reduction gearbox whose efficiency is 95 percent. What is the output torque?  


1
Expert's answer
2021-11-01T12:53:47-0400

The torque developed by the motor:


"T=P/\omega=P/(2\pi n)."

The output torque due to the reducer is


"T_r=rT=rP/(2\pi n)."


The total output torque:

"T_o=rT=\eta rP/(2\pi n),\\
T_o=0.95·4(30·10^3)/(2\pi(3600/60))=302\text{ N}·\text{m}."